what volume of concentrated HCL that is 36% HCL by mass and has density of 1.18 g/cm3 is required yo produce 8.25L of a solution with pH of 1.77

Grade:12th Pass

2 Answers

Sunil Kumar FP
9 years ago

ph=1.77
-log[H+]=1.77
[H+]=.588M
mole of HCl=.588*8.25=.485mole
let the volume of Hcl be V
now .36*V*1.18=.485 V=1.14L
TEJKIRAT SINGH
9 years ago
ph=1.77-log[H+]=1.77[H+]=.588Mmole of HCl=.588*8.25=.485molelet the volume of Hcl be Vnow .36*V*1.18=.485 V=1.14L